OpenStudy (anonymous):

let f(x)=x^2-x-6 find x when f(x)=6

OpenStudy (lgbasallote):

i'll do the first step for you...f(x) = 6 means... x^2 - x - 6 = 6 what do you think comes next?

OpenStudy (anonymous):

x^2-x-6-6=0

OpenStudy (anonymous):

subtract 6?

OpenStudy (anonymous):

x^2-x=0

OpenStudy (lgbasallote):

right @moriyah.claxton it's x^2 - x - 12 = 0 @beketso

OpenStudy (anonymous):

ur right.igba

OpenStudy (anonymous):

i'm not sure what to do next

OpenStudy (anonymous):

(x-4)(x+3)=0 (factorise)

OpenStudy (anonymous):

x=4 OR x=-3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

pleasure